# A Self-Explainable Heterogeneous GNN for Relational Deep Learning

### v1

@@ A.4 Synthetic datasets
Table 6: $F_{1}$ metric with standard deviations for synthetic datasets
A.5 Real world setting
@@
Medical database with 31 tables (node types) [^3] https://eicu-crd.mit.edu. The task is predicting the duration of a patient’s stay in the eICU after admission, modeled as binary node classification by thresholding duration at 20 hours to achieve two balanced classes. To create clusters of nodes, where each cluster is represented by a single new node that replaces all the nodes within that cluster, we utilized a categorical attribute for each table that is best suited for clustering the specific table.
@@
Database  [^4] https://relational-data.org/dataset/Mondial containing data from multiple geographical web data sources [50]. We predict the religion of a country as Christian (positive) with 114 instances vs. all other religions with 71 instances. In this dataset, clustering of tables is done using DBSCAN [51] clustering algorithm.
@@
A.6 Real world setting results
EICU MONDIAL ErgastF1 rel-f1-dnf rel-f1-top3
MLP 0.53(±${\scriptstyle(\pm 0.02)}$ 0.52(±${\scriptstyle(\pm 0.00)}$ 0.50(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$ 0.48(±${\scriptstyle(\pm 0.00)}$
GCN 0.89(±${\scriptstyle(\pm 0.00)}$ 0.60(±${\scriptstyle(\pm 0.02)}$ 0.50(±${\scriptstyle(\pm 0.01)}$ 0.57(±${\scriptstyle(\pm 0.01)}$ 0.52(±${\scriptstyle(\pm 0.02)}$
RGCN 0.70(±${\scriptstyle(\pm 0.00)}$ 0.53(±${\scriptstyle(\pm 0.08)}$ 0.57(±${\scriptstyle(\pm 0.01)}$ 0.44(±${\scriptstyle(\pm 0.02)}$ 0.54(±${\scriptstyle(\pm 0.02)}$
R-HGNN 0.61(±${\scriptstyle(\pm 0.00)}$ 0.61(±${\scriptstyle(\pm 0.01)}$ 0.72(±${\scriptstyle(\pm 0.02)}$ 0.60(±${\scriptstyle(\pm 0.01)}$ 0.63(±${\scriptstyle(\pm 0.02)}$
HGN 0.75(±${\scriptstyle(\pm 0.00)}$ 0.72(±${\scriptstyle(\pm 0.01)}$ 0.70(±${\scriptstyle(\pm 0.04)}$ 0.61(±${\scriptstyle(\pm 0.02)}$ 0.61(±${\scriptstyle(\pm 0.01)}$
GTN 0.56(±${\scriptstyle(\pm 0.02)}$ 0.38(±${\scriptstyle(\pm 0.01)}$ 0.60(±${\scriptstyle(\pm 0.01)}$ 0.41(±${\scriptstyle(\pm 0.02)}$ 0.45(±${\scriptstyle(\pm 0.01)}$
Fast-GTN 0.46(±${\scriptstyle(\pm 0.03)}$ 0.39(±${\scriptstyle(\pm 0.04)}$ 0.60(±${\scriptstyle(\pm 0.03)}$ 0.51(±${\scriptstyle(\pm 0.01)}$ 0.50(±${\scriptstyle(\pm 0.02)}$
MP-GNN 0.87(±${\scriptstyle(\pm 0.02)}$ 0.36(±${\scriptstyle(\pm 0.06)}$ 0.71(±${\scriptstyle(\pm 0.01)}$ 0.54(±${\scriptstyle(\pm 0.02)}$ 0.52(±${\scriptstyle(\pm 0.02)}$
MPS-GNN 0.92(±${\scriptstyle(\pm 0.01)}$ 0.74(±${\scriptstyle(\pm 0.01)}$ 0.83(±${\scriptstyle(\pm 0.02)}$ 0.62(±${\scriptstyle(\pm 0.02)}$ 0.65(±${\scriptstyle(\pm 0.01)}$
A.7 Necessity calculation
To assess whether an explanation is faithful to the model, it is crucial to identify the elements that significantly influence the model’s performance. In other words, if any part of this explanation is removed, the prediction should deteriorate. This concept is known as Necessity, as it evaluates which components of the explanation are indispensable for accurate predictions. To quantify this, a distance metric is used, measuring the difference in predicted probabilities between the original predictions and those obtained after masking parts of the explanation (i.e. deleting some instances of meta-paths). This metric is calculated as:
$\textsc{Nec}=\frac{1}{N}{\sum}_{v=1}^{N}(p_{v}(\mathcal{G})-p_{v}(\mathcal{G}^{\prime}))$ (8)
where $v$ is a target node, $\mathcal{G}$ is the original graph and $p_{v}$ denotes the probability associated with the predicted class. $\mathcal{G^{\prime}}$ is obtained by removing certain meta-path occurrences and $p_{v}(\mathcal{G^{\prime}})$ is the probability associated at the class predicted with $p_{v}$. Since MPS-GNN utilizes just the explanation for making prediction one should expect that removing some part of the explanation, has as effect a decrease in prediction $F_{1}$. Figure 5 shows five plots, one for each dataset. As in Table 3 each graph illustrates the effects of removing 25%, 50%, and 75% of the meta-path occurrences in terms of changes in $F_{1}$ and necessity between original and modified graphs [49].
Figure 5: Changes to $F_{1}$ and posterior probability difference (necessity) when removing 25%, 50%, and 75% of the learned meta-path occurrences for the real-world datasets.
A.8 Number of parameters
@@ A.9 Execution times
A.10 Scoring Function Lookahead Illustration
@@
A.11 Temporal experiment
Recently, a novel benchmark, rel-bench [31], has been introduced. This benchmark consists of multiple relational databases represented as temporal graphs. Additionally, the authors propose RDL, a temporal-aware relational message-passing model. It’s important to note that the released temporal graphs are not directly compatible with static models. Therefore, in this section, we reconstruct one dataset from the rel-bench repository in a static format and apply our method to this static representation of the graph (setting in table 14). The tasks are: (1) dnf, predicting whether a driver will fail to finish a race within the next month, and (2) top3, predicting if a driver will place in the top 3. To handle these temporal tasks, we treated each instance of a node at different timestamps as distinct nodes with separate label predictions. Table 15 reports the $F_{1}$-scores for MPS-GNN and the baselines. Our approach outperforms all competitors, including RDL which performs lower due to overpredicting the majority class. Note that RDL cannot be straightforwardly applied to our other datasets, being designed for temporal datasets. In Section A.11.1, we provide the necessity calculations for these datasets.
@@ A.11 Temporal experiment
Table 14: Setting of temporal real-world dataset.
@@
A.11.1 Necessity calculation in temporal tasks
In this section, we present the necessity calculation for the temporal datasets. Specifically, as detailed in Appendix A.7, Table 16 and figure6 demonstrates that removing certain identified meta-paths results in decreased performance ($F_{1}$) and an increased necessity value. This finding confirms that the learned meta-paths are essential for the prediction task in these datasets as well.
@@ A.11.1 Necessity calculation in temporal tasks
A.12 Non-GNN models on real-world databases
For the MONDIAL and ErgastF1 databases, non-GNN methods have shown competitive performance in the past. For example, [52, 53] report $F_{1}$ scores of $0.78$ and $0.77$ on MONDIAL, $0.4$ and $0.3$ points higher than MPS-GNN. However, these results are achieved on a simplified version of the database with only 12 tables, requiring manual feature selection. In contrast, MPS-GNN is applied directly to the raw input data across all 40 tables. The non-GNN methods use Multi-relational Bayes Net Classifiers and Simple Decision Forests, where reducing the number of tables and relations aids performance. By comparison, MPS-GNN is designed to handle scenarios with a large number of relations effectively.

### v2

@@ 5.2 Q2: MPS-GNN surpasses competitors in real world databases, learning relevant meta-paths
Our approach is particularly useful for predictive tasks in relational databases with multiple tables, where features for a target entity may involve statistics from related tables. To address the second research question, we thus focused on three relational databases with many tables: EICU, a medical database with 31 tables, where we predict patient stay duration in the eICU, modeled as binary node classification by thresholding duration at 20 hours to achieve two balanced classes.; MONDIAL, a geographic database where the task is predicting whether a country’s religion is Christian; and ErgastF1, containing Formula 1 data, where the task is predicting the winner of a race in a binary classification task where target nodes are represented by a combination of race and pilot. The databases were transformed into graphs as explained in Section 3; for disconnected components, we enhanced connectivity by clustering rows of auxiliary tables. Additional details for the datasets and the procedure are in the Appendix A.3.
Results
EICU MONDIAL ErgastF1
MLP 0.53(±${\scriptstyle(\pm 0.02)}$ 0.52(±${\scriptstyle(\pm 0.00)}$ 0.50(±${\scriptstyle(\pm 0.00)}$
GCN 0.89(±${\scriptstyle(\pm 0.00)}$ 0.60(±${\scriptstyle(\pm 0.02)}$ 0.50(±${\scriptstyle(\pm 0.01)}$
RGCN 0.70(±${\scriptstyle(\pm 0.00)}$ 0.53(±${\scriptstyle(\pm 0.08)}$ 0.57(±${\scriptstyle(\pm 0.01)}$
R-HGNN 0.61(±${\scriptstyle(\pm 0.00)}$ 0.61(±${\scriptstyle(\pm 0.01)}$ 0.72(±${\scriptstyle(\pm 0.02)}$
HGN 0.75(±${\scriptstyle(\pm 0.00)}$ 0.72(±${\scriptstyle(\pm 0.01)}$ 0.70(±${\scriptstyle(\pm 0.04)}$
GTN 0.56(±${\scriptstyle(\pm 0.02)}$ 0.38(±${\scriptstyle(\pm 0.01)}$ 0.60(±${\scriptstyle(\pm 0.01)}$
Fast-GTN 0.46(±${\scriptstyle(\pm 0.03)}$ 0.39(±${\scriptstyle(\pm 0.04)}$ 0.60(±${\scriptstyle(\pm 0.03)}$
MP-GNN 0.87(±${\scriptstyle(\pm 0.02)}$ 0.36(±${\scriptstyle(\pm 0.06)}$ 0.71(±${\scriptstyle(\pm 0.01)}$
MPS-GNN 0.92(±${\scriptstyle(\pm 0.01)}$ 0.74(±${\scriptstyle(\pm 0.01)}$ 0.83(±${\scriptstyle(\pm 0.02)}$
Table 2 presents the $F_{1}$ scores of MPS-GNN and its competitors across three real-world databases, averaged over 5 runs with different seeds. The poor performance of MLP clearly indicates that using target node features only is insufficient for classification. Plain GCN, which treats the graph as homogeneous, performs well only on the EICU dataset, where node degree differences exist between positive and negative nodes. Heterogeneous GNN methods also struggle with these datasets, especially MONDIAL, where most approaches fail to outperform a simple MLP, and only one (HGN) manages to substantially outperform the non-heterogeneous baseline (GCN). MP-GNN does not provide the performance boost that was observed when applied to knowledge graphs (Ferrini et al., 2024), confirming our intuition that existential quantification of meta-path is insufficient when dealing with relational databases. On the other hand, MPS-GNN manages to substantially outperform all competitors, thanks to its ability to identify meta-paths that are informative thanks to the statistics that can be computed over their realizations, as shown in the following. It is worth noting that this result is achieved with one/two orders of magnitude fewer parameters than the runner-ups, namely HGN and R-HGNN. See Table 9 in the Appendix for the details. Additionally, Table 10 in the Appendix shows that MPS-GNN has competitive execution times with respect to other heterogeneous GNN approaches, thanks to its ability to focus training on relevant meta-path induced subgraphs.
Identified Meta-Paths
Figure 7: Extracted meta-paths for the three real world datasets.
Finally, in Appendix A.8 we present an experimental evaluation where MPS-GNN is adapted to deal with temporal databases and tasks (Robinson et al., 2024), showing how it outperforms its competitors also in this context. Summing up, these results enable us to confidently answer Q2 in the affirmative.
5.3 Q3: MPS-GNN is a self-explainable method
To address the third research question, we assessed the faithfulness of the extracted meta-paths. The meta-paths identified by the model from the graph are evaluated based on the complementary metrics of sufficiency and necessity. High sufficiency implies that changing the complement to the explanation (leaving the explanation unchanged) should not affect the model’s output. High necessity implies that altering the explanation itself (leaving the complement unchanged) should result in a change in the model’s output. It is easy to show that our approach is inherently sufficient. Indeed, the computational graph of MPS-GNN consists solely of the subgraph containing the occurrences of the identified meta-paths. Necessity, on the other hand, is calculated as a distance metric, measuring the difference in predicted probabilities between the original predictions and those obtained after masking parts of the explanation (i.e. deleting some instances of meta-paths). Defined as $\textsc{Nec}=\frac{1}{N}{\sum}_{v=1}^{N}(p_{v}(\mathcal{G})-p_{v}(\mathcal{G}^{\prime}))$ where $v$ is a target node, $\mathcal{G}$ is the original graph and $p_{v}$ denotes the probability associated with the predicted class. $\mathcal{G^{\prime}}$ is obtained by removing certain meta-path occurrences (i.e. randomly removing some branches of the identified meta-paths) and $p_{v}(\mathcal{G^{\prime}})$ is the probability associated at the class predicted with $p_{v}$. Since MPS-GNN utilizes just the explanation for making prediction one should expect that removing some part of the explanation, has as effect a decrease in prediction $F_{1}$.
Figure 8: Changes to $F_{1}$ and posterior probability difference (necessity) when removing 25%, 50%, and 75% of the learned meta-path occurrences for the real-world datasets with MPS-GNN, dashed line, and MP-GNN, solid line.
Figure 8 illustrates the effects of removing 25%, 50%, and 75% of the meta-path occurrences in terms of changes in $F_{1}$ and necessity between original and modified graphs (Azzolin et al., 2025). In addition to the results for MPS-GNN (dashed line), the figure includes the results for MP-GNN (solid line), which is also a self-explainable GNN model according to the reasoning in Section 4.5. In all datasets, there’s a noticeable decline in $F_{1}$ performance for MPS-GNN and a steep increase in probability difference, suggesting that the learned meta-paths are also necessary. These results clearly indicate the faithfulness of the explanations of MPS-GNN. MP-GNN has a similar behaviour, albeit with lower $F_{1}$ with respect to MPS-GNN because of its lower expressivity. The only exception is the MONDIAL dataset, where MP-GNN fails to learn any relevant pattern, resulting in a very low $F_{1}$ score that remains approximately constant when removing meta-path occurrences.
6 Conclusion
We introduced a novel approach to identify relevant meta-paths of relations for node classification tasks in heterogeneous graphs with a potentially large number of different relations, notably graphs derived from relational databases. Compared to earlier work, our approach does not require user supervision and learns meta-paths for predictive features defined by aggregate statistics over meta-path occurrences. The explainability of our method is particularly beneficial for sensitive domains like medical or financial data, where it helps address fairness concerns by providing insights into predictions. Experiments demonstrate advantages in accuracy and explainability.
Limitations and future works
At this stage, our method is tailored for binary node classification but can be extended to multiclass classification using standard one-vs-all strategies. However, this approach is inefficient, as it requires repeating the entire process for each class. The underlying principles of our local greedy meta-path construction based on scoring potential informativeness of relations directly applies also in the multiclass case. What would be needed in a multiclass adaptation of the approach is to replace the scalar functions (1),(2), and scalar node weights $w_{u}$, with vector-valued versions, and to modify the loss function (3) accordingly. Similarly, node regression is currently not supported and would require a modification to the way relations are scored. Another limitation of the scoring function is its reliance on a well-connected graph structure; when target nodes have neighbors that are not connected to other target nodes, the scoring function requires a preprocessing step to create supernodes within the neighborhood, as implemented in certain real-world scenarios.
An interesting direction for future work involves incorporating temporal information into the process of learning optimal meta-paths. By doing so, the model would be capable of accounting for the temporal dimension, enabling it to better capture time-dependent relationships and dynamics within the graph.
7 Acknowledgments
Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. Grant Agreement no. 101120763 - TANGO. AL acknowledges the support of the MUR PNRR project FAIR - Future AI Research (PE00000013) funded by the NextGenerationEU.
Appendix A Appendix
A.1 Second iteration loss computation
In this Section we show the loss computation and minimization for relation "d" of the example in Figure 5. Note that $Z$ refers to the learned parameters related to the features of the target nodes in the previous iteration with relation $b$ (Figure 3(b)) and $\begin{bmatrix}1\\ 0\end{bmatrix}$ and $\begin{bmatrix}0\\ 1\end{bmatrix}$ represent the features of $prescription$ nodes.
$Z$ $\gg 0$
$F(\mathcal{B}_{2},d)$ $=Z\Theta^{T}\begin{bmatrix}0\\ 1\end{bmatrix}+2Z\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}(w_{8}+w_{9})+Z\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}(w_{10}+w_{11})$
$F(\mathcal{B}_{3},d)$ $=2Z\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}(w_{8}+w_{9})+Z\Theta^{T}\begin{bmatrix}0\\ 1\end{bmatrix}(w_{10}+w_{11})+Z\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}(w_{12})$
$L(d)$ $=\min_{\Theta,w}\sigma\left(F(\mathcal{B}_{3},d)-F(\mathcal{B}_{2},d)\right)$
$=\min_{\Theta,w}\sigma\left(Z\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}(2w_{8}+2w_{9}+w_{12}-2w_{8}-2w_{9}-w_{10}-w_{11})\right.$
$\quad\left.+Z\Theta^{T}\begin{bmatrix}0\\ 1\end{bmatrix}(w_{10}+w_{11}-1)\right)$
$=\min_{\Theta,w}\sigma\left(Z\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}(w_{12}-w_{10}-w_{11})+Z\Theta^{T}\begin{bmatrix}0\\ 1\end{bmatrix}(w_{10}+w_{11}-1)\right)$
$\Theta^{T}\begin{bmatrix}1\\ 0\end{bmatrix}$ $\gg 0;\quad\Theta^{T}\begin{bmatrix}0\\ 1\end{bmatrix}=0;\quad w_{12}=0;\quad w_{10},w_{11}=1$
A.2 Ablation study
In this section, we highlight the importance of the skip connection $W_{1}^{(l)}h_{v}^{(0)}$ considered in equation 7. Table 3 and 4 demonstrate respectively the significance of skip connections in the synthetic and real-world settings. Without considering the initial target node features, the $F_{1}$ score drops significantly, underscoring the critical role these features play.
S1 S2 S3 S4 S5 S6 S7 S8
[…]
EICU MONDIAL ErgastF1
[…]
A.3 Real world setting
@@
Medical database with 31 tables (node types) [^4] https://eicu-crd.mit.edu from Johnson et al. (2021). The task is predicting the duration of a patient’s stay in the eICU after admission, modeled as binary node classification by thresholding duration at 20 hours to achieve two balanced classes. To create clusters of nodes, where each cluster is represented by a single new node that replaces all the nodes within that cluster, we utilized a categorical attribute for each table that is best suited for clustering the specific table.
@@
Database  [^5] https://relational-data.org/dataset/Mondial containing data from multiple geographical web data sources (May, 1999). We predict the religion of a country as Christian (positive) with 114 instances vs. all other religions with 71 instances. In this dataset, clustering of tables is done using DBSCAN (Ester et al., 1996) clustering algorithm.
@@
A.4 Details of competitors’ architectures
In Table 8, we present the hyperparameters for all competitors and MPS-GNN on the real-world datasets. For the synthetic cases, the only difference lies in the # layers, which is adjusted to match the length of the correct meta-paths for the competitors. For the losses CE is cross-entropy and nll is negative log likelihood.
Table 8: Hyperparameters of competitors and MPS-GNN for the real-world datasets. The optimizer is omitted from the table as it is Adam for all models. lr denotes the learning rate, wd represents the weight decay, and Patience indicates the early stopping patience (if applicable).
Hyperparameters # layers Embedding dim. lr wd # epochs Patience Loss
MLP 2 32 0.01 0.0005 500 50 nll
GCN 2 32 0.01 0.0005 500 50 nll
RGCN 2 32 0.01 0.0005 500 50 nll
R-HGNN 2 64 0.001 0 200 50 CE
HGN 2 64 0.0001 0.0005 300 30 nll
GTN 2 64 0.01 0.001 200 - CE
Fast-GTN 2 64 0.01 0.001 200 - CE
MP-GNN 2 32 0.01 0.0005 500 50 nll
MPS-GNN 2 32 0.01 0.0005 500 50 nll
A.5 Number of parameters
@@ A.6 Execution times
A.7 Scoring Function Lookahead Illustration
@@
A.8 Temporal experiment
Recently, a novel benchmark, rel-bench Robinson et al. (2024), has been introduced. This benchmark consists of multiple relational databases represented as temporal graphs. Additionally, the authors propose RDL, a temporal-aware relational message-passing model. It’s important to note that the released temporal graphs are not directly compatible with static models. Therefore, in this section, we reconstruct one dataset from the rel-bench repository in a static format and apply our method to this static representation of the graph (setting in table 12). The tasks are: (1) dnf, predicting whether a driver will fail to finish a race within the next month, and (2) top3, predicting if a driver will place in the top 3. To handle these temporal tasks, we treated each instance of a node at different timestamps as distinct nodes with separate label predictions. Table 13 reports the $F_{1}$-scores for MPS-GNN and the baselines. Our approach outperforms all competitors, including RDL which performs lower due to overpredicting the majority class. Note that RDL cannot be straightforwardly applied to our other datasets, being designed for temporal datasets. In Section A.8.1, we provide the necessity calculations for these datasets.
Table 12: Setting of temporal real-world dataset.
@@ A.8 Temporal experiment
A.8.1 Necessity calculation in temporal tasks
In this section, we present the necessity calculation for the temporal datasets. Specifically, as detailed in Section 5.3, Table 14 and figure9 demonstrates that removing certain identified meta-paths results in decreased performance ($F_{1}$) and an increased necessity value. This finding confirms that the learned meta-paths are essential for the prediction task in these datasets as well.
@@ A.8.1 Necessity calculation in temporal tasks
A.9 Non-GNN models on real-world databases
For the MONDIAL and ErgastF1 databases, non-GNN methods have shown competitive performance in the past. For example, Schulte et al. (2013); Bina et al. (2013) report $F_{1}$ scores of $0.78$ and $0.77$ on MONDIAL, $0.4$ and $0.3$ points higher than MPS-GNN. However, these results are achieved on a simplified version of the database with only 12 tables, requiring manual feature selection. In contrast, MPS-GNN is applied directly to the raw input data across all 40 tables. The non-GNN methods use Multi-relational Bayes Net Classifiers and Simple Decision Forests, where reducing the number of tables and relations aids performance. By comparison, MPS-GNN is designed to handle scenarios with a large number of relations effectively.
A.10 Toy example with more complex features
Figure 10: Toy example similar to Figure 1 with more complex features on prescription nodes. As depicted in the legend (right), the first two positions of the feature vector determine an exempt prescription ($[1,0]$) or a not exempt prescription ($[0,1]$). The last value determine the price of the prescription in dollar. The highlighted subgraph shows a prototypical counts-of-counts pattern characterising positive patients, namely having exempt prescriptions where the total cost is more then 100 dollars, each containing at least two medications
In Figure 10, we present a scenario similar to the one depicted in Figure 1. The key difference lies in the prescription nodes, which now have more complex feature representations. Specifically, the last value of the feature vector corresponds to the price of each prescription.
The first iteration of the scoring function follows the same process as in the previous toy example, as illustrated in Figures 3(a) and 3(b). The most notable aspect, however, is the second iteration, where the target nodes are prescription nodes, similar to Figure 5, where the model evaluates relation $d$ and  $c$. Here, we provide the computations involved in minimizing the loss. Notably, only relation $d$ allows the loss to be minimized to 0. Therefore, relation $d$ will be the one included in the identified meta-path
Minimization relation $c$
$Z$ $\gg 0$
$F(\mathcal{B}_{2},c)$ $=2Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 60\end{bmatrix}+Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{13})+Z\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}$
$F(\mathcal{B}_{3},c)$ $=2Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 60\end{bmatrix}+Z\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}+Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{13})$
$L(c)$ $=\min_{\Theta,w}\sigma\left(F(\mathcal{B}_{3},d)-F(\mathcal{B}_{2},d)\right)$
$=\min_{\Theta,w}\sigma\left(0\right)$
$=\frac{1}{2}$
Minimization relation $d$
$Z$ $\gg 0$
$F(\mathcal{B}_{2},d)$ $=2Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 60\end{bmatrix}(w_{8}+w_{9})+Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{10}+w_{11})+Z\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}$
$F(\mathcal{B}_{3},d)$ $=2Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 60\end{bmatrix}(w_{8}+w_{9})+Z\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}(w_{10}+w_{11})+Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{12})$
$L(d)$ $=\min_{\Theta,w}\sigma\left(F(\mathcal{B}_{3},d)-F(\mathcal{B}_{2},d)\right)$
$=\min_{\Theta,w}\sigma\left(Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{12})+Z\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}(w_{10}+w_{11})-Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{10}+w_{11})-Z\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}\right)$
$=\min_{\Theta,w}\sigma\left(Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{12})+Z\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}(w_{10}+w_{11}-1)-Z\Theta^{T}\begin{bmatrix}1\\ 0\\ 70\end{bmatrix}(w_{10}+w_{11})\right)$
$\Theta^{T}\begin{bmatrix}0\\ 1\\ 20\end{bmatrix}$ $\gg 0;\quad w_{10},w_{11},w_{12}=0$